OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y? x = 8, y = 4 x = 10, y = 2 x = 18, y = -6 x = 20, y = -8 Answer?

Mehek (mehek14):

which one do you think?

OpenStudy (anonymous):

idk

Mehek (mehek14):

ok so adding any of the 2 numbers, the answer is 12 so now look at the second part of your question

Mehek (mehek14):

The difference of x and two times y is 6. What are the values of x and y?

Mehek (mehek14):

Meaning we look at the answer choices and subtract them

Mehek (mehek14):

look at A 8 - 2*4 = ?

Mehek (mehek14):

8 - 8 = 12??? I don't think so...

OpenStudy (anonymous):

sorry i was rushing to fast

Mehek (mehek14):

it's fine

Mehek (mehek14):

so does 8-8 = 6?

OpenStudy (anonymous):

no

Mehek (mehek14):

so A isn't the answer

Mehek (mehek14):

Moving on to B 10 - 2*2 = 6?

OpenStudy (anonymous):

yes

Mehek (mehek14):

so that's the answer

OpenStudy (anonymous):

thanks :)

Mehek (mehek14):

yw \(\ddot\smile\)